OpenStudy (kellyspeakslouder):

I know this is a lot, but I can't put the equations in the right form with this computer. f(x) = 3x + 2; g(x) = 3x - 5 Find f/g. A. (f/g)(x) = Quantity three x minus five divided by three x plus two.; domain {x|x ≠ - Two over three. } B. (f/g)(x) = Quantity three x plus two divided by three x minus five.; domain {x|x ≠ Five over three. } C. (f/g)(x) = Quantity three x minus five divided by three x plus two; domain {x|x ≠ Five over three. } D. (f/g)(x) = Quantity three x plus two divided by three x minus five; domain {x|x ≠ - Two over three. }

OpenStudy (kellyspeakslouder):

@jim_thompson5910 I have this question then one more after this. I know this is a mess.

OpenStudy (danjs):

similar thing, have to divide them \[\large (f/g)(x) = \frac{ f(x) }{ g(x) } = \frac{ 3x + 2 }{ 3x-5 }\]

jimthompson5910 (jim_thompson5910):

are you able to post a screenshot?

OpenStudy (kellyspeakslouder):

no :( I am sorry.. and I'll divide

jimthompson5910 (jim_thompson5910):

@DanJS is correct. You just need to now find the domain. Set the denominator equal to zero and then solve for x. That will show you which value is going to be excluded from the domain.

OpenStudy (kellyspeakslouder):

Idk I'm not getting any of the answer options.

jimthompson5910 (jim_thompson5910):

3x-5 = 0 solve for x to get x = ??

OpenStudy (kellyspeakslouder):

2?

jimthompson5910 (jim_thompson5910):

add 5 to both sides first

jimthompson5910 (jim_thompson5910):

then divide both sides by 3

OpenStudy (kellyspeakslouder):

-2?

jimthompson5910 (jim_thompson5910):

3x-5 = 0 3x-5+5 = 0+5 3x = 5 x = ??? (it's not -2)

OpenStudy (kellyspeakslouder):

would the answer be B or C?

jimthompson5910 (jim_thompson5910):

3x = 5 3x/3 = 5/3 ... divide both sides by 3 x = 5/3

jimthompson5910 (jim_thompson5910):

To avoid division by zero errors, x = 5/3 must be excluded out of the domain

OpenStudy (kellyspeakslouder):

that's where I was horribly confused

OpenStudy (kellyspeakslouder):

C?

jimthompson5910 (jim_thompson5910):

look carefully at what @DanJS posted

OpenStudy (kellyspeakslouder):

ohhh B

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (kellyspeakslouder):

thank you!

jimthompson5910 (jim_thompson5910):

no problem